# A Liquid-Core Fiber Platform for Classical and Entangled Two-Photon Absorption Measurements

### v1

@@ 4 Results and Discussion
SPDC loss % 99.96a 24
Intact pairs % 0.00002a 58
E2PA cross section UB cm2 $(4.0\pm 1.6)\times 10^{-21}$ $(2.1\pm 0.5)\times 10^{-25}$
$R^{\mathrm{UB}}$ 17
@@
We can compare the upper bound set here to the upper bound set in Ref. 16 for the same sample as shown in Table 1. Since the E2PA cross section is known to scale with parameters of the excitation beam that varied between the two experiments, we cannot compare the two numbers alone. We instead compare the upper bounds using help from the probabilistic model (see for example Ref. 16) and arrive at the ratio $R^{\mathrm{UB}}$,
@@
Although we do not know the exact entanglement areas of the SPDC used in either experiment, the same ppKTP crystal, pump laser and pump focusing conditions were used in both experiments. In the best-case scenario, if the spatial correlations of the SPDC are maintained as it propagates through the optical system the entanglement area should scale with the beam size at the image plane of the crystal. Since the optical system is designed to image the crystal into the sample, we can use this logic to estimate the ratio of the entanglement areas in the previous experiment to that in the current experiment as the ratio of the respective beam sizes of the SPDC at the sample position. Conveniently, the upper bounds on the entanglement areas are set by these beam sizes. In free space, the E2PA cross section upper bound was set to $(2.1\pm 0.5)\times 10^{-25}$ cm2 for $T_{e}=1620$ fs and $A_{e}$ in the range $2.1$ $\mu$m2 to $13,700$ $\mu$m2. Plugging in the numbers we arrive at $R^{\mathrm{UB}}$ of 17. This result indicates that the current experiment sets an upper bound that is 17-fold larger than the previous experiment, and is thus 17-fold less stringent.
5 Conclusions
We presented a toluene-filled fiber platform for two-photon excited fluorescence measurements. We used this platform to measure classical two-photon absorption of various concentrations of the sample AF455. The results of the C2PA measurements show that 2PA can be measured at extremely low powers ($1.75$ nW) using an excitation volume of only 7.3 nL. This power is 45-fold lower than we achieved previously in free-space under similar conditions (Ref. 16). Furthermore, we derived a classical two-photon absorption cross section for AF455 of (390$\pm$80) GM. These results emphasize the advantages of liquid-core-fiber platforms for sensing applications where one could instead use higher powers to sense very low concentrations of analytes.
With photon pair excitation, we saw no evidence of a signal, and were ultimately limited by low coupling efficiency of the spatially multi-mode SPDC into fiber. Despite this high coupling loss, our calculation of $R^{\mathrm{UB}}$ in Section 4, which uses $\sigma_{E}^{\mathrm{UB}}$ to account for many experimental factors such as loss and the broadening of the entanglement time, shows that the bound we set on the entangled two-photon absorption cross section is 17-fold larger than that set in Ref. 16. Thus, this technique does not provide a higher E2PA sensitivity compared to the free-space technique. However, since the SPDC loss reduces the upper bound in a quadratic manner, efforts to decrease the loss would greatly improve the sensitivity of this technique. This work furthers the growing body of research that has presented null results when attempting to measure entangled two-photon absorption (Refs. 21, 16, 22, 23, 46, 24, 25, 26, 27, 28, 29, 30).
@@ 5 Conclusions
This work was supported by NIST and by the NSF Physics Frontier Center at JILA (PHY 2317149) and by the NSF-STROBE center (DMR 1548924). We thank S. Mukherjee, A. McLean, H. Greene, K. Thatcher, J. Sipe and C. Drago for valuable suggestions and discussions on the experimental design. We are grateful to T. Loon-Seng Tan and T. Cooper for providing molecular samples. Certain commercial equipment, instruments, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identification is not intended to imply recommendation or endorsement by NIST, nor is it intended to imply that the materials or equipment identified are necessarily the best available for the purpose.
{suppinfo}
Additional information on experimental characterization, data acquisition and data analysis.
6 Experimental Characterization
In this section we discuss the experimental characterizations we performed to derive C2PA cross sections using Eq. (9) and an E2PA cross section upper bound using Eq. (10). First, we discuss the transmission efficiency of the fiber. Then, we discuss the characterization of propagation losses of the fiber by imaging fiber scatter. Afterwards, we discuss images of the excitation and fluorescence light at the fiber face for qualitative identification of the mode content of the fiber. Finally, we discuss our characterization of the setup’s dispersion.
Before C2PA experiments we maximize the transmission of the 810 nm laser light through the fiber. We measure the power before and after the fiber using power meters, and the ratio of the power transmitted out $W_{\mathrm{out}}$ (W) to the power sent in $W_{\mathrm{in}}$ (W) is what we call our transmission efficiency,
@@ 6 Experimental Characterization
This efficiency is a product of coupling ($\eta_{C}$), absorption ($\eta_{A}$) and scattering ($\eta_{S}$) efficiencies. The latter two efficiencies are evaluated at the excitation wavelength ($\lambda_{e}=810$ nm) and for the entire length of the fiber $l$ (cm). The best alignment resulted in $\eta_{T}>55\%$. However, during our measurements $\eta_{T}\approx 43\%$ but varied from data series to data series. We followed the same procedure to measure SPDC transmission efficiency, however we use an sCMOS camera instead of a power meter to measure the incident and output photon count rate. The highest value achieved was $\eta_{T}\approx 0.058\%$.
We characterize the loss of light along the length $z$ of the fiber by imaging the scattering. The intensity of the scattering is proportional to the power $W(z)$, and thus should exponentially decay according to
@@
With only the solvent in the fiber ($c=0$), we send a 458 nm (near the peak of our sample’s emission spectrum) alignment laser through the fiber and image the scattering on a smart phone camera using long exposure settings as shown in Fig. 7. The image is integrated along the width of the fiber (vertically) to find an integrated intensity as a function of fiber length. The intensity is fit to the exponential in Eq. (4) to derive $a_{\mathrm{sol}}(458\,\mathrm{nm})+\mu(458\,\mathrm{nm})=0.093$ cm-1 for fiber 1 (Fig. 7(a)) and $a_{\mathrm{sol}}(458\,\mathrm{nm})+\mu(458\,\mathrm{nm})=0.034$ cm-1 for fiber 2 (Fig. 7(b)). For 458 nm fluorescence, traveling the entire length of the fiber corresponds to a loss of 96.8% or 70.6% of the light for fibers 1 and 2 respectively. We take the same measurement with 750 nm (a visible wavelength close to our excitation wavelength) light and observe no evidence of decay of the scatter along the length of the fiber. We estimate that $a_{\mathrm{sol}}(750\,\mathrm{nm})+\mu(750\,\mathrm{nm})$ is not large enough to measure this way. As mentioned in Section 2, the values of $a_{\mathrm{sol}}(\lambda)$ are known. From literature,39 $a_{\mathrm{sol}}(750\,$nm)=0.0036)=0.0036) = 0.0036 cm-1 and $a_{\mathrm{sol}}(810\,$nm)=0.0030)=0.0030) = 0.0030 cm-1. If we set $\mu(750\,\mathrm{nm})=\mu(810\,\mathrm{nm})=0$, we find that these absorption coefficients correspond to a loss of 12.5% and 10.5%, respectively, of the light along the length of the fiber. Considering with our best alignment $\eta_{T}>55$%, and that some fraction of that light is lost due to the coupling efficiency in addition to the 10.5% absorption loss, we estimate $\mu(810\,\mathrm{nm})\approx 0$. This low scattering loss is consistent with the light primarily occupying the fundamental mode of the fiber. From here we can derive that for $\eta_{T}=43\%$, $\eta_{C}=48\%$.
@@
For SPDC, we are not able to measure $\eta_{S}$ due to the low photon rate, and thus cannot quantify the division of losses between $\eta_{C}$ and $\eta_{S}$. However, we can consider that $\eta_{S}$ will likely only differ from that for laser light if the mode content in the fiber is different. If the SPDC does occupy more higher-order modes than laser light, the higher-order modes tend to have high $\eta_{S}$. This would cause a quick depletion of the modes from the fiber leading to minimal contributions to an E2PA signal. Furthermore, higher-order modes have spatial profiles that are less ideal for 2PA. Thus without losing the bulk of the proper physical model, we will group the scattering losses of the higher-order modes into the coupling loss. Then we can use the same $\eta_{S}$ as that determined for the fundamental mode measured with laser light, and the coupling efficiency will serve as an effective fundamental mode coupling efficiency ($\mu(810\,\mathrm{nm})\approx 0$, $\eta_{C}=0.065\%$). We assume that the losses are equivalent for both the signal and idler modes that are coupled into fiber.
@@
The fluorescence is collected at the front of the fiber using the lens optimized to focus 810 nm light into the fiber, which roughly collimates the visible fluorescence. The light is reflected at the dichroic beamsplitter, and another lens focuses the image of the fiber face onto the EMCCD, as shown in Fig. 2(d). In this image, the high intensity bright spot is from fluorescence guided through the core. Surrounding that spot is darkness likely from the cladding of the fiber, indicating that there is little light propagating through the cladding of the fiber. Radially outward from the cladding, light forms an offset circular shape and is likely from light guided through the core of the tubing sleeve as can be seen by comparison to the digital microscope image in Fig. 2(b). A larger circular ring of light surrounds that core, and is likely from light guided through the “cladding” of the tubing sleeve. Thus, the spatial mode content of the fluorescence in the fiber is more complex than the simple calculation in Section 2 assumes, however the highest intensity mode appears to be the fundamental mode of the fiber.
To measure the free-space group delay dispersion (GDD), $D_{0}$ (fs2), accumulated by the laser pulse as it propagates through the optical setup, we use the GDD tuning function of the laser. We increase the dispersion compensation in increments of 500 fs2 and take C2PA measurements at each step. The step with the maximum C2PA signal corresponds to the value at which the GDD is optimally compensated at the input of the fiber. We measure $D_{0}\approx 2000$ fs2, $D_{0}\approx 2000$ fs2 and $D_{0}\approx 4000$ fs2 for experiments 1, 2 and 3 respectively. For experiments 2 and 3 the free-space dispersion is compensated for using the internal tuning of the laser system. To estimate $D_{0}$ accumulated by the SPDC as it propagates through the optical setup, we calculate the GDD of each optical element in its path from the center of the crystal to the entrance of the fiber. We estimate $D_{0}\approx 2100$ fs2. To determine the fiber group velocity dispersion (GVD) $\beta$ (fs2 cm-1), a COMSOL simulation is used to model the filled fiber. The low-order modes are solved for. For the fundamental mode, $\beta$ is identical to $\beta$ of toluene, thus there are no contributions from waveguide dispersion. However, for higher-order modes it is likely that $\beta$ is affected by waveguide dispersion. These higher-order modes are not considered in the calculations of Section 9 or Section 10, and $\beta$ of the fundamental mode is used.
7 Data Acquisition
@@
8 Data normalization
In this section we discuss the data normalization used to account for fluctuations in excitation power. For this we made use of $W_{\mathrm{out}}$, which is measured for each frame, since it accounts for potential changes in $\eta_{C}$ due to laser pointing drifts unlike $W_{\mathrm{in}}$.
@@ 8 Data normalization
For the SPDC measurement, a quadrant detector is used to monitor the pump power while the sCMOS camera is used to measure $W_{\mathrm{out}}$. The output power varied over a range of 24% of the maximum value, and decreased monotonically relative to the pump power by 24% over the course of the 24 hour measurement. The latter value is directly related to the decrease in coupling efficiency. The value $\eta_{C}=0.065\%$ is the average coupling efficiency.
9 Calculating a C2PA Cross Section
In this section we discuss the assumptions and equations used to calculate a C2PEF signal. We arrive at an equation to derive a C2PA cross section. Finally, we use these equations to model the concentration dependence of the signal and compare with our results.
@@ 9 Calculating a C2PA Cross Section
Here we define $\kappa(\lambda)$ as the fraction of fluorescence that can be collected by the fiber and directed out in the direction of the detector. We assume that the fluorescence is emitted isotropically and that the fraction of that light collected can be described by the solid angle of a cone with apex angle $2(90^{\circ}-\theta_{c})$, where $\theta_{c}$ is the critical angle (Section 2). Then we can write,
@@
$=$ $\eta_{A}(\lambda,z)\eta_{S}(\lambda,z)\gamma_{0}(\lambda),$
where $\eta_{A}(\lambda,z)$ and $\eta_{S}(\lambda,z)$ are the absorption and scattering efficiencies in fiber (Section 6), $\mathcal{T}(\lambda)$ and $\mathcal{R}(\lambda)$ are the transmittance and reflectance of an optic (w = window, l = lens, d = dichroic beamsplitter and f = spectral filter), and $\mathrm{QE}(\lambda)$ (cnt photon-1) is the quantum efficiency of the camera.
$N_{C}(z)=\frac{1}{2}\sigma_{C}n\times$ (12)
[…]
where $\sigma_{C}$ (1 GM = 10-50 cm4 s photon-1 fluorophore-1) is the C2PA cross section, $n$ (fluorophores cm-3) is the number density of the fluorophores and $\phi(x,y,z,t)$ (photons cm-2 s-1) is the photon flux of the laser beam. The factor of $1/2$ carries units of excitations per photons absorbed. The temporal and transverse spatial profiles of the laser beam are approximated by Gaussian distributions. The transverse spatial profile of the light inside the fiber will differ from a Gaussian distribution if the light occupies the higher-order modes of the fiber, but in this calculation we make a few approximations based on the assumption that all the light is in the fundamental mode (Section 6). Assuming the laser is always on, $\phi(x,y,z,t)$ takes the form
where $\phi_{0}(z)$ (photons cm-2 s-1) is the peak photon flux as a function of $z$, $\tau(z)$ (fs) is the FWHM pulse duration and $d_{0}$ (cm) is the FWHM beam width. We use COMSOL to solve for $d_{0}$ by first solving for the effective mode area $A_{\textrm{eff}}$ (cm2) of the fundamental mode47 and solving
$d_{0}=\sqrt{\frac{2\textrm{ln}2A_{\textrm{eff}}}{\pi}}.$ (14)
[…]
$=$ $\frac{W(z)}{h\nu},$
We can solve for $\phi_{0}(z)$ by inputting Eq. (9) into Eq. (15) and integrating over $x$, $y$ and $t$,
Now we can rewrite Eq. (9) using these equations as
[…]
where $\lambda_{e}=810$ nm.
Then we can solve for the C2PA cross section,
@@
All the parameters in Eq. (9) are known through experiments or simulations. The parameters $d_{0}$ and $\beta$ are estimated using COMSOL Multiphysics Simulation software, $F_{C}/W_{0}^{2}$ (cnt s-1 $\mu$W-2) is the fit to our experimental C2PEF data (and $W_{0}$ is estimated using measured $W(l)$), $n$ is measured using a spectrophotometer, $a_{\mathrm{sol}}(\lambda)$ is from literature39, $\mu(\lambda)$ is determined from scattering measurements (Section 6), $\tau_{0}$ is measured using a SwampOptics Grenouille 8-50-USB, $D_{0}$ is measured, $k(\lambda)$ is calculated, $\Phi(\lambda)$ is known from published AF455 measurements44, 45 and $\gamma(z,\lambda)$ is calculated based on optics’ specifications, our scattering measurements and the published extinction coefficient and spectra of AF45545. Table 2 and Table 3 summarize parameter values used in this calculation and that described in Section 10.

### v2

@@ Methods
SPDC loss % 73a 24
Intact pairs % 25a 58
E2PA cross-section UB cm2 $(5.8\pm 2.3)\times 10^{-24}$ $(2.1\pm 0.5)\times 10^{-25}$
$R^{\mathrm{UB}}$ 8.5
@@
Although E2PEF was indistinguishable from zero, there are numerous conclusions to be drawn from this measurement. The C2PEF dataset for the $2.30$ mM sample is plotted alongside our E2PEF measurement in Fig. 5(b). Here we show measured fluorescence count rates as a function of the average power of either the SPDC (orange) or laser (blue) source. For the E2PEF measurement, we couple about 36.5 pW ($\approx 1.9$ photons pulse-1) of SPDC into the fiber. This plot demonstrates the difference between the SPDC power and the minimum laser power, which are only 48-fold apart.
We can compare these E2PA results with those done in free space (Table 1). Notably, the power of SPDC at the sample is 60-fold lower in the present experiment. This is because the free-space experiment uses all the photons from $\approx 740$ spatial modes, whereas the fiber acts as a few-mode filter. Despite this lower power, the peak photon flux is $7.7\times 10^{19}$ photons cm-2s-1. In the free-space studies the SPDC peak photon flux was $2.1\times 10^{18}$ photons cm-2s-1, which is 37-fold lower. It should be noted that some of the advantage of the higher SPDC photon flux is hindered by the lower fraction of intact photon pairs propagating in the fiber, which is 25% for fiber and 58% for free space. The intact pair rate is accounted for using the estimated Klyshko efficiency in the calculations in the Supporting Information.
Using Eq. (Calculating an E2PA Cross-Section Upper Bound) in the Supporting Information, we set an upper bound on the E2PA cross-section of this sample at $(5.8\pm 2.3)\times 10^{-24}$ cm2. This upper bound is for an entanglement time of 1070 fs and an entanglement area in the range $2.1$ $\mu$m2 to $18$ $\mu$m2. This entanglement time is an estimated value for the SPDC at the entrance of the fiber adapter. The broadening of the entanglement time in the fiber is among numerous parameters accounted for in the calculation of the upper bound (see Supporting Information).
We can compare the upper bound set here to the upper bound set in Ref. 16 for the same sample as shown in Table 1. Since the E2PA cross-section is known to scale with parameters of the excitation beam that varied between the two experiments, we cannot compare the two numbers alone. For instance, in an experiment where the SPDC beam is tightly focused at the sample, the entanglement area typically decreases proportionally, leading to a larger E2PA cross-section compared to experiments with gentler focusing. Similary, if the SPDC is sent through dispersive elements, the temporal separation between photons in a pair broadens—especially for SPDC sources with broad bandwidths—resulting in an increased entanglement time and a reduced E2PA cross-section. To properly compare the upper bounds, we use the probabilistic model (see for example Ref. 16) and arrive at the ratio $R^{\mathrm{UB}}$,
@@
Although we do not know the exact entanglement areas of the SPDC used in either experiment, the same ppKTP crystal, pump laser and pump focusing conditions were used in both experiments. Assuming that the spatial correlations of the SPDC are maintained as it propagates through the optical system, the entanglement area should scale with the beam size at the image plane of the crystal. Since the optical system is designed to image the crystal into the sample, we can use this logic to estimate the ratio of the entanglement areas in the previous experiment to that in the current experiment as the ratio of the respective beam sizes of the SPDC at the sample position. For a fair comparison, we use the beam size of all $\approx 740$ spatial modes at the focus of the lens used for fiber coupling, which is $6,350$ $\mu$m2. In free space, the beam size at its focus in the sample was $13,700$ $\mu$m2. Plugging in the numbers we arrive at $R^{\mathrm{UB}}$ of 8.5. This result indicates that the current experiment sets an upper bound that is 8.5-fold larger than the previous experiment, and is thus 8.5-fold less stringent.
Conclusions
We presented a toluene-filled fiber platform for two-photon excited fluorescence measurements. We used this platform to measure classical two-photon absorption of various concentrations of the sample AF455. The results of the C2PA measurements show that 2PA can be measured at extremely low powers ($1.75$ nW) using an excitation volume of only 7.3 nL. This power is 45-fold lower than we achieved previously in free-space under similar conditions (Ref. 16). Furthermore, we derived a classical two-photon absorption cross-section for AF455 of (390$\pm$80) GM. These results emphasize the advantages of liquid-core-fiber platforms for sensing applications where one could instead use higher powers to sense very low concentrations of analytes.
With photon pair excitation, we saw no evidence of a signal, and set an upper bound on the E2PA cross-section of AF455. Our calculation of $R^{\mathrm{UB}}$, which uses $\sigma_{E}^{\mathrm{UB}}$ to account for many experimental factors such as loss and the broadening of the entanglement time, shows that the bound we set on the entangled two-photon absorption cross-section is 8.5-fold larger than that set in Ref. 16. Thus, this technique does not provide a higher E2PA sensitivity compared to the free-space technique. However, since the SPDC loss reduces the upper bound in a quadratic manner, efforts to decrease the loss would greatly improve the sensitivity of this technique. This work furthers the growing body of research that has presented null results when attempting to measure entangled two-photon absorption (Refs. 23, 16, 24, 25, 26, 27, 28, 29, 30, 31, 32, 22).
@@ Conclusions
This work was supported by NIST and by the NSF Physics Frontier Center at JILA (PHY 2317149) and by the NSF-STROBE center (DMR 1548924). We thank L. K. Shalm, S. Mukherjee, A. McLean, H. Greene, K. Thatcher, J. Sipe and C. Drago for valuable suggestions and discussions on the experimental design. We are grateful to T. Loon-Seng Tan and T. Cooper for providing molecular samples. Certain commercial equipment, instruments, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identification is not intended to imply recommendation or endorsement by NIST, nor is it intended to imply that the materials or equipment identified are necessarily the best available for the purpose.
Supporting Information
Experimental Characterization
In this section we discuss the experimental characterizations we performed to derive C2PA cross-sections using Eq. (Calculating a C2PA Cross-Section) and an E2PA cross-section upper bound using Eq. (Calculating an E2PA Cross-Section Upper Bound). First, we discuss the transmission efficiency and propagation losses of the fiber. Next, we discuss the considerations needed to switch from single-mode laser light to multimode SPDC. We characterize the number of spatial modes and the Klyshko efficiency. Afterwards, we discuss images of the excitation and fluorescence light at the fiber face for qualitative identification of the mode content of the fiber. Finally, we discuss our characterization of the setup’s dispersion.
Before C2PA experiments we maximize the transmission of the 810 nm laser light through the fiber. We measure the power before and after the fiber using power meters, and the ratio of the power transmitted out $W_{\mathrm{out}}$ (W) to the incident power $W_{\mathrm{in}}$ (W) is what we call our transmission efficiency,
@@
This efficiency is a product of coupling ($\eta_{C}$), absorption ($\eta_{A}$) and scattering ($\eta_{S}$) efficiencies. The latter two efficiencies are evaluated at the excitation wavelength ($\lambda_{e}=810$ nm) and for the entire length of the fiber $l$ (cm). The best alignment resulted in $\eta_{T}>55\%$. However, during our measurements $\eta_{T}\approx 43\%$ but varied from data series to data series. This value for $\eta_{T}$ serves as a best estimate for the transmission efficiency for any aligned single spatial mode source at the same frequency, including a single mode of SPDC. Similarly, the values of the components of $\eta_{T}$: $\eta_{C}$, $\eta_{A}$ and $\eta_{S}$, derived using this measurement and the scattering measurements as described below, also serve as a best estimate of these parameters for any single mode of the same frequency.
We characterize the loss of light along the length $z$ of the fiber by imaging the scattering. The intensity of the scattering is proportional to the power $W(z)$ (W), and thus should exponentially decay according to
@@
With only the solvent in the fiber ($c=0$), we send a 458 nm (near the peak of our sample’s emission spectrum) alignment laser through the fiber and image the scattering on a smart phone camera using long exposure settings as shown in Fig. 7. The image is integrated along the width of the fiber (vertically) to find an integrated intensity as a function of fiber length. The intensity is fit to the exponential in Eq. (4) to derive $a_{\mathrm{sol}}(458\,\mathrm{nm})+\mu(458\,\mathrm{nm})=0.093$ cm-1 for fiber 1 (Fig. 7(a)) and $a_{\mathrm{sol}}(458\,\mathrm{nm})+\mu(458\,\mathrm{nm})=0.034$ cm-1 for fiber 2 (Fig. 7(b)). For 458 nm fluorescence, traveling the entire length of the fiber corresponds to a loss of 96.8% or 70.6% of the light for fibers 1 and 2 respectively. We take the same measurement with 750 nm (a visible wavelength close to our excitation wavelength) light and observe no evidence of decay of the scatter along the length of the fiber. We estimate that $a_{\mathrm{sol}}(750\,\mathrm{nm})+\mu(750\,\mathrm{nm})$ is not large enough to measure this way. As mentioned in the main text, the values of $a_{\mathrm{sol}}(\lambda)$ are known. From literature,40 $a_{\mathrm{sol}}(750\,$nm)=0.0036)=0.0036) = 0.0036 cm-1 and $a_{\mathrm{sol}}(810\,$nm)=0.0030)=0.0030) = 0.0030 cm-1. If we set $\mu(750\,\mathrm{nm})=\mu(810\,\mathrm{nm})=0$, we find that these absorption coefficients correspond to a loss of 12.5% and 10.5%, respectively, of the light along the length of the fiber. Considering with our best alignment $\eta_{T}>55$%, and that some fraction of that light is lost due to the coupling efficiency in addition to the 10.5% absorption loss, we estimate $\mu(810\,\mathrm{nm})\approx 0$. This low scattering loss is consistent with the light primarily occupying the fundamental mode of the fiber. From here we can derive that for $\eta_{T}=43\%$, $\eta_{C}=48\%$.
@@
For SPDC, we are not able to measure $\eta_{S}$ due to the low photon rate, and thus cannot quantify the division of losses between $\eta_{C}$ and $\eta_{S}$. However, we can consider that $\eta_{C}$ and $\eta_{S}$ will likely only differ from that for laser light if the mode content in the fiber is different. If the SPDC does occupy more higher-order modes than laser light, the higher-order modes tend to have high $\eta_{S}$. This would cause a quick depletion of the modes from the fiber leading to minimal contributions to an E2PA signal. Furthermore, higher-order modes have spatial profiles that are less ideal for 2PA. Thus without losing the bulk of the proper physical model, we will operate under the assumption that only the fundamental mode is relevant in the E2PA measurement.
We followed a similar procedure to that used for measuring the transmission efficiency of the laser in order to estimate the number of SPDC spatial modes. Here we write the transmission efficiency as,
$\eta_{T}$ $=$ $\frac{Q_{\mathrm{out}}M}{Q_{\mathrm{in}}^{\mathrm{mm}}}=\frac{Q_{\mathrm{out}}}{Q_{\mathrm{in}}}$
where $M$ is the total number of spatial modes of the incident light, $Q_{\mathrm{in}}$ (photons s-1) and $Q_{\mathrm{out}}$ (photons s-1) are the incident and output photon count rate and $Q_{\mathrm{in}}^{\mathrm{mm}}$ (photons s-1) is the incident photon count rates measured for a multimode (mm) light source. Unless specified with the labeling “mm”, all other quantities are for a single spatial mode. The values of $Q_{\mathrm{out}}$ and $Q_{\mathrm{in}}^{\mathrm{mm}}$ can be measured directly using an sCMOS camera. The incident photon count rate is an estimate for the incident photon count rate of the single mode coupled into fiber. The highest value achieved for the ratio of output photon count rate to incident multimode photon count rate is $Q_{\mathrm{out}}/Q_{\mathrm{in}}^{\mathrm{mm}}\approx 0.058\%$. We can plug in the values of $\eta_{S}$, $\eta_{A}$ and $\eta_{C}$ determined using a laser source, as well as $Q_{\mathrm{out}}/Q_{\mathrm{in}}^{\mathrm{mm}}$ into Eq. (Experimental Characterization) to estimate $M\approx 740$. This estimate for the number of spatial modes serves as a lower bound since up to 16 modes can be coupled into fiber as calculated using Eq. (1) in the main text.
Thus far we have discussed coupling efficiency $\eta_{C}$ defined as the ratio of single photons incident on the fiber to single photons output from the fiber—all of which originate from the same single mode. This is in contrast to the Klyshko efficiency44—a quantity typically measured for photon pair sources—defined as the probability of detecting a photon conditioned on the successful detection of its partner photon. For this work, we use a theoretical value for the Klyshko efficiency to aid in our estimation of intact photon pair rate in the fiber, $Q_{\mathrm{pairs}}$ (photon pairs s-1). We input various parameters of our pump beam, crystal, lenses and fiber into SPDCalc45 which calculates the overlap integral of three Gaussian spatial modes—one for signal photons, one for idler photons, and one for the collected single mode in fiber—along the length of the crystal and estimates $\eta^{{}^{\prime}}_{K}=0.94$. This differs from a measured Klyshko efficiency, $\eta_{K}$, because it does not account for any single photon loss between photon pair generation and collection into fiber. We can relate the two by,
$\eta_{K}$ $=$ $\eta^{{}^{\prime}}_{K}\eta_{F}\eta_{C}$
$=$ $\frac{2Q_{\mathrm{pairs}}(z=0)}{Q_{\mathrm{singles}}(z=0)},$
where $\eta_{F}$ is the free-space transmission efficiency between the center of the crystal to the fiber and $Q_{\mathrm{pairs}}(z)$ and $Q_{\mathrm{singles}}(z)$ (photons s-1) are the rates of photon pairs and single photons, respectively, in fiber. In later sections, we will refer to $Q_{\mathrm{singles}}(z)$ simply as $Q(z)$. The factor of 2 is used to align with the standard definition of $\eta_{K}$, which is typically measured using two fibers. In writing this equation and Eq. (Experimental Characterization), we have assumed that the efficiencies of the various frequency modes are equal. We find that $\eta_{K}=0.25$.
@@
The fluorescence is collected at the front of the fiber using the lens optimized to focus 810 nm light into the fiber, which roughly collimates the visible fluorescence. The light is reflected at the dichroic beamsplitter, and another lens focuses the image of the fiber face onto the EMCCD, as shown in Fig. 2(d). In this image, the high intensity bright spot is from fluorescence guided through the core. Surrounding that spot is darkness likely from the cladding of the fiber, indicating that there is little light propagating through the cladding of the fiber. Radially outward from the cladding, light forms an offset circular shape and is likely from light guided through the core of the tubing sleeve as can be seen by comparison to the digital microscope image in Fig. 2(b). A larger circular ring of light surrounds that core, and is likely from light guided through the “cladding” of the tubing sleeve. Thus, the spatial mode content of the fluorescence in the fiber is more complex than Eq. (1) in the main text assumes, however the highest intensity mode appears to be the fundamental mode of the fiber.
To measure the free-space group delay dispersion (GDD), $D_{0}$ (fs2), accumulated by the laser pulse as it propagates through the optical setup, we use the GDD tuning function of the laser. We increase the dispersion compensation in increments of 500 fs2 and take C2PA measurements at each step. The step with the maximum C2PA signal corresponds to the value at which the GDD is optimally compensated at the input of the fiber. We measure $D_{0}\approx 2000$ fs2, $D_{0}\approx 2000$ fs2 and $D_{0}\approx 4000$ fs2 for experiments 1, 2 and 3 respectively. For experiments 2 and 3 the free-space dispersion is compensated for using the internal tuning of the laser system. To estimate $D_{0}$ accumulated by the SPDC as it propagates through the optical setup, we calculate the GDD of each optical element in its path from the center of the crystal to the entrance of the fiber. We estimate $D_{0}\approx 2100$ fs2. To determine the fiber group velocity dispersion (GVD) $\beta$ (fs2 cm-1), a COMSOL simulation is used to model the filled fiber. The low-order modes are solved for. For the fundamental mode, $\beta$ is identical to $\beta$ of toluene, thus there are no contributions from waveguide dispersion. However, for higher-order modes it is likely that $\beta$ is affected by waveguide dispersion. These higher-order modes are not considered in the calculations of the C2PA cross-section and E2PA cross-section upper bound, and $\beta$ of the fundamental mode is used.
Data Acquisition
@@
Data Normalization
In this section we discuss the data normalization used to account for fluctuations in excitation power. For this, we made use of $W_{\mathrm{out}}$, which is measured for each frame, since it accounts for potential changes in $\eta_{C}$ due to laser pointing drifts unlike $W_{\mathrm{in}}$.
@@
For the SPDC measurement, a quadrant detector is used to monitor the pump power while the sCMOS camera is used to measure $W_{\mathrm{out}}$. The output power varied over a range of 24% of the maximum value, and decreased monotonically relative to the pump power by 24% over the course of the 24 hour measurement. The latter value is directly related to a decrease in coupling efficiency. The value $Q_{\mathrm{out}}/Q_{\mathrm{in}}^{\mathrm{mm}}=0.058\%$ is the average value over the course of the measurement. Similarly, $\eta_{C}$, which is measured using the laser, is the average from three datasets each measured over the course of more than 24 hours and was thus also subject to drifts in coupling efficiency.
Calculating a C2PA Cross-Section
In this section we discuss the assumptions and equations used to calculate a C2PEF signal. We arrive at an equation to derive a C2PA cross-section. Finally, we use these equations to model the concentration dependence of the signal and compare with our results.
@@
Here we define $\kappa(\lambda)$ as the fraction of fluorescence that can be collected by the fiber and directed out in the direction of the detector. We assume that the fluorescence is emitted isotropically and that the fraction of that light collected can be described by the solid angle of a cone with apex angle $2(90^{\circ}-\theta_{c})$, where $\theta_{c}$ is the critical angle (see main text). Then we can write,
@@
$=$ $\eta_{A}(\lambda,z)\eta_{S}(\lambda,z)\gamma_{0}(\lambda),$
where $\eta_{A}(\lambda,z)$ and $\eta_{S}(\lambda,z)$ are the absorption and scattering efficiencies in fiber (as described earlier), $\mathcal{T}(\lambda)$ and $\mathcal{R}(\lambda)$ are the transmittance and reflectance of an optic (w = window, l = lens, d = dichroic beamsplitter and f = spectral filter), and $\mathrm{QE}(\lambda)$ (cnt photon-1) is the quantum efficiency of the camera.
$N_{C}(z)=\frac{1}{2}\sigma_{C}n\times$ (14)
[…]
where $\sigma_{C}$ (1 GM = 10-50 cm4 s photon-1 fluorophore-1) is the C2PA cross-section, $n$ (fluorophores cm-3) is the number density of the fluorophores and $\phi(x,y,z,t)$ (photons cm-2 s-1) is the photon flux of the laser beam. The factor of $1/2$ carries units of excitations per photons absorbed. The temporal and transverse spatial profiles of the laser beam are approximated by Gaussian distributions. The transverse spatial profile of the light inside the fiber will differ from a Gaussian distribution if the light occupies the higher-order modes of the fiber, but in this calculation we make a few approximations based on the assumption that all the light is in the fundamental mode. The spatial integrals extend from negative infinity to positive infinity and are equivalent to integrals that extend over only the core of the fiber. Assuming the laser is always on, $\phi(x,y,z,t)$ takes the form
where $\phi_{0}(z)$ (photons cm-2 s-1) is the peak photon flux as a function of $z$, $\tau(z)$ (fs) is the FWHM pulse duration and $d_{0}$ (cm) is the FWHM beam width. We use COMSOL to solve for $d_{0}$ by first solving for the effective mode area $A_{\textrm{eff}}$ (cm2) of the fundamental mode49 and solving
$d_{0}=\sqrt{\frac{2\textrm{ln}2A_{\textrm{eff}}}{\pi}}.$ (16)
[…]
$=$ $\frac{W(z)}{h\nu},$